b) A family of 2 adults and 3 kids go and watch a movie. The price of an adult ticket is m and the price of a child ticket is n such that:
2m + 3n = 28 and n = 12 - m
I thought I would solve it by simply doing
2m+3(0)=28 and then solve for n the same but this is not working out.
2m=28
m=14


Answer by rapaljer(4671)

You are assuming that n=0, which is not the case. What you really need to do is use the n=12-m to substitute into the first equation in place of the n like this:

n=12-m
2m+3n=28

2m+3( ) = 28
2m+3(12-m)=28

Now, solve for m:
2m + 36-3m = 28
36-m=28
-m=28-36
-m=-8
m=8

Now, substitute m=8 into the n equation:
n=12-m
n=12-8
n=4

Dr. Rapalje